OpenStudy (anonymous):

indicated the result when b is added to the numerator and subtracted from the denominator of the fraction a over c?

OpenStudy (callisto):

fraction a over c \[=> \frac{a}{c}\] b is added to the numerator a \[=> \frac{a+b}{c}\] b is ... subtracted from the denominator \[=> \frac{a+b}{c-b}\]